OpenStudy (deoxna):

A particle undergoes SHM. The graph below shows how the kinetic energy EK of the particle varies with the displacement x of the particle from equilibrium.

OpenStudy (deoxna):

I'll upload the word document, but I need help with question 6.

OpenStudy (anonymous):

can you save that as a screen shot or something?

OpenStudy (deoxna):

This the relevant graph, but my question is the following, how to find the maximum velocity using the maximum kinetic energy?

OpenStudy (anonymous):

Two points probably... ie \[KE _{1} / KE _{2}\]

OpenStudy (anonymous):

mass cancels... hmm still doesn't give enough info. though... do you know anything about the spring (k for example)?

OpenStudy (deoxna):

I calculated the spring using the formula ME=PE+KE=(1/2)(K)(A)², where the amplitude is 0.05 m PE+KE=0.06=1/2k(0.05)² -> k=48

OpenStudy (anonymous):

good point... let me think about it for a minute or two

OpenStudy (anonymous):

still there?

OpenStudy (deoxna):

yes, I'm going over my notes to find other formulas...

OpenStudy (anonymous):

Think I came up with something...

OpenStudy (anonymous):

\[V _{\max} = (k/m)*x _{\max}\]

OpenStudy (anonymous):

you can use that in \[1/2*m*(V _{\max})^{2} = .06J\]

OpenStudy (anonymous):

v= sq root (k/m) x

OpenStudy (anonymous):

Yes wrong bracket use in the equation editor sorry..

OpenStudy (deoxna):

I found the equation: \[v=\sqrt{\frac{ k }{ m }}\sqrt{A^2-x²}\], which simplifies to V=wA

OpenStudy (anonymous):

yes...it's right but how u'll get omega?

OpenStudy (anonymous):

in fact which is ur next question...find d angular frequency

OpenStudy (anonymous):

there is another relation also k=m (w)^2...but it also won't help.

OpenStudy (deoxna):

Also, -kx=ma

OpenStudy (anonymous):

but i dont think there is any use of acceleration term in this question.

OpenStudy (anonymous):

from where u got this question...some assignment?

OpenStudy (deoxna):

Yeah, but what I'm thinking is that this might be unsolvable without mass, and my teacher may have forgotten to include it

OpenStudy (anonymous):

Yes..it seems like that.

OpenStudy (deoxna):

Or, we have been looking at spring-mass systems, whereas pendulums are also shm that dont depend on mass.

OpenStudy (anonymous):

but u need length of the pendulum

OpenStudy (deoxna):

yeah... I just realized that...

OpenStudy (deoxna):

Anyway, I'll just let it be and tell him I couldn't solve it. But anyway thanks alot for helping me.

OpenStudy (anonymous):

:)